A dehtdrated patients needs a 3.82% saline IV. Unfortunately, the hospital has bags of 4% and 3% saline solutios. How much liters of each of these solutios should be mixed together to yield 5 liters of the desired oncentration?
Answer by reviewermath(1029)

0.04x + 0.03(5-x) = 0.0382(5)
0.01x = 0.041
x = 4.1
5-x = 0.9
Answer: 4.1 liters of 4% saline solution and 0.9 liters of 3% saline solution.